what is the flight time of a cannon fires a cannonball with a muzzle velocity of 50.0 m/sec at an angle of 40 degrees.


1
Expert's answer
2021-05-04T11:30:49-0400

To be given in question

Initial velocity (u)=50m/sec

"g=9.8m\/sec^2"

"\theta=40°"

To be asked in question

Total time in flight

T=?

We know that

"T=\frac{2usin\theta}{g}\rightarrow(1)"

Put value in eqution (1)

"T=\frac{2\times 50\times sin(40°)}{9.8}"

T=6.55 sec